# Gabor systems with Hermite functions of order $n$ and oversampling greater than $n+1$ which are not frames

### v1

@@ 3. Results
Due to the results of Gröchenig and Lyubarskii [6][7] we know that if we have a Gabor system over a lattice $\Lambda$ with ${\rm vol}({\mathbb{R}}^{2}/\Lambda)<1/(n+1)$, then $\mathcal{G}(h_{n},\Lambda)$ is a frame. ∎
Note, that this does not exclude the option that $\mathcal{Z}h_{n}$ has more than $n+1$ zeros in $[0,1)^{2}$.
Corollary 3.4.
There exist Gabor systems with Hermite functions of order $n\geq 1$ over a periodic index set of density strictly greater than $n+1$ which are not frames.
The proof of Corollary 3.4 is by example. We start with the first Hermite function.
Example 1. Consider the Zak transform of $h_{1}$. Since $h_{1}$ is odd, we know that
@@
In particular, the Zak transform $\mathcal{Z}h_{1}$ has (at least) 3 zeros in $[0,1)^{2}$. Thus, there exists a periodic configuration $\Gamma$ of density 3 such that $\mathcal{G}(h_{1},\Gamma)$ is not a frame. As we know the exact location of the 3 zeros, we know $\Gamma$, which we depict in Figure 2. Concretely, the Gabor system $\mathcal{G}\left(h_{1},{\mathbb{Z}}^{2}\cup(({\mathbb{Z}}+\frac{1}{2})\times{\mathbb{Z}})\cup({\mathbb{Z}}\times({\mathbb{Z}}+\frac{1}{2}))\right)$ is not a frame.
Figure 2. The set $\Gamma=\bigcup_{m=1}^{3}({\mathbb{Z}}^{2}+z_{m})$, $z_{m}\in\{(0,0),(\frac{1}{2},0),(0,\frac{1}{2})\}$ and its decomposition into 3 relatively shifted copies of the integer lattice ${\mathbb{Z}}^{2}$.
Example 2 We consider the third Hermite function $h_{3}$. In [10], Lemvig lists zeros of $\mathcal{Z}\mathcal{D}_{\sqrt{a}}^{-1}h_{4\ell+3}$, $\ell\in{\mathbb{N}}_{0}$, $a\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},2,3,4\}$. These were earlier discovered by Boon, Zak, and Zucker [1] in a different context for $\mathcal{Z}_{a}h_{4\ell+3}$. Due to a different normalization of the Fourier transform $\mathcal{Z}_{a}$ is defined slightly different in [1] compared to (3.12). We have the following known zeros (further zeros may exist), as listed in [1] (subject to our normalization):
@@
Note that the first three zeros in any of the above lists are due to the parity of $h_{4\ell+3}$. The cases for $\frac{1}{\sqrt{a}}$ are covered by an application of the Poisson summation formula, similarly to (2.28), which results in simply rotating the rectangle in Figure 3 by 90 degrees.

### v2

@@ 3. Results
Due to the results of Gröchenig and Lyubarskii [15][16] we know that a Gabor system $\mathcal{G}(h_{n},\Lambda)$ over a lattice $\Lambda$ with $D(\Lambda)>(n+1)$ is a frame. If $\{z_{1},\ldots,z_{M}\}$, $M>n+1$, are zeros of $\mathcal{Z}h_{n}$ and $\Gamma=\Gamma({\mathbb{Z}}^{2},\{z_{m}\}_{m=1}^{M})$, then $\Gamma$ has density $D(\Gamma)=M>n+1$. As $\mathcal{G}(h_{n},\Gamma)$ is not a frame by Theorem 3.1, $\Gamma$ cannot be a lattice. ∎
Note that this does not exclude the option that $\mathcal{Z}h_{n}$ has more than $n+1$ zeros in $[0,1)^{2}$. This is a key message of the paper, which leads to the following observation:
There exist Gabor systems with Hermite functions of order $n\geq 1$ over periodic index sets of density strictly greater than $n+1$ which are not frames.
Although the following result is an easy consequence of Theorem 3.1, it marks the starting point of the study of Gabor systems $\mathcal{G}(h_{n},\Gamma)$ with Hermite functions, $n\geq 1$, where $\Gamma$ is not a lattice but a (more general) periodic set.
Theorem 3.4.
Set $z_{1}=(0,0)$, $z_{2}=(1/2,0)$, and $z_{3}=(0,1/2)$. Then, the periodic configuration $\Gamma({\mathbb{Z}}^{2},\{z_{m}\}_{m=1}^{3})$ has density 3 and the Gabor system $\mathcal{G}\left(h_{1},\Gamma({\mathbb{Z}}^{2},\{z_{m}\}_{m=1}^{3})\right)$ is not a Gabor frame for $L^{2}({\mathbb{R}})$.
Proof.
Since $h_{1}(t)=-h_{1}(-t)$ is an odd function, we know that
@@
The result now follows from Theorem 3.1. ∎
This provides the negative answer as announced in the introduction. We exploited the fact that the Zak transform $\mathcal{Z}h_{1}$ has (at least) three zeros in $[0,1)^{2}$. Thus, by Theorem 3.1, there exists a periodic configuration $\Gamma$ of density 3 such that $\mathcal{G}(h_{1},\Gamma)$ is not a frame. In addition, we know the exact location of the three zeros making $\Gamma$ explicit, which we depict in Figure 2. While it is not a lattice, it still exhibits periodicities and symmetries.
Figure 2. The set $\Gamma=\bigcup_{m=1}^{3}({\mathbb{Z}}^{2}+z_{m})$, $z_{m}\in\{(0,0),(\frac{1}{2},0),(0,\frac{1}{2})\}$ and its decomposition into three relatively shifted copies of the integer lattice ${\mathbb{Z}}^{2}$.
Examples. We now give further examples of Gabor systems which are not frames and where the density exceeds the order of the Hermite function by more than 1. Consider the third Hermite function $h_{3}$. In [21], Lemvig lists zeros of $\mathcal{Z}\mathcal{D}_{\sqrt{a}}^{-1}h_{4\ell+3}$, $\ell\in{\mathbb{N}}_{0}$, $a\in\{\frac{1}{4},\frac{1}{3},\frac{1}{2},2,3,4\}$. Boon, Zak, and Zucker earlier discovered these in a different context for $\mathcal{Z}_{\sqrt{a}}h_{4\ell+3}$ [3]. Due to a different normalization of the Fourier transform, $\mathcal{Z}_{a}$ is defined slightly differently in [3] compared to (3.13). We have the following known zeros (further zeros may exist), as listed in [3] (subject to our normalization):
@@
Note that, after dilation, the four zeros with $\omega=0$ can be used to construct examples of a semi-regular lattice $\Gamma_{N}\times{\mathbb{Z}}$ of density $N\in\{2,3,4\}$ such that the Gabor system with $\mathcal{D}_{\sqrt{3}}^{-1}h_{3}$ is not a frame. Denote by $z_{1}=(0,0)$, $z_{2}=\left(\frac{1}{3},0\right)$, $z_{3}=\left(\frac{1}{2},0\right)$, and $z_{4}=\left(\frac{2}{3},0\right)$, then
(3.19) $\mathcal{G}\left(\mathcal{D}_{\sqrt{3}}^{-1}h_{3},\Gamma_{N}\times{\mathbb{Z}}\right),\ \Gamma_{N}=\bigcup_{k=1}^{N}({\mathbb{Z}}+z_{k})\quad\text{ is not aframe for }N=2,3,4.$
@@
The location of the zeros in $[0,2)\times[0,\frac{1}{2})$ is depicted in Figure 5.
@@
Note that the first three zeros in any of the above lists are due to the parity of $h_{4\ell+3}$. The cases for $\frac{1}{\sqrt{a}}$ are covered by an application of the Poisson summation formula, similarly to (2.28), which results in simply rotating the rectangle in figure 3, 4, and 5 by 90 degrees.
@@
4. Open questions and new zeros
While our results give new insights into the mysterious world of Gabor systems with Hermite functions, they immediately raise more questions.
  1. (Q1)
    For any $n\geq 1$, is there a $\Gamma$ with $D(\Gamma)>n+1$ such that $\mathcal{G}(h_{n},\Gamma)$ is not a frame?
  2. (Q2)
    How many zeros does the Zak transform of $h_{n}$ have?
  3. (Q3)
    How do the zeros of $\mathcal{Z}_{a}h_{n}$ differ from the zeros of $\mathcal{Z}h_{n}$?
In general, it seems that considering $\mathcal{Z}_{a}h_{n}$, $a\in{\mathbb{R}}_{+}\setminus\{1\}$, introduces new non-trivial zeros for $n\geq 2$, as shown in [3], [19], [21].
Conjectures about the frame property of Gabor systems with Hermite functions tend to be risky. However, we believe that the following statement about the Zak transform holds.
Conjecture 4.1.
For a generalized first Hermite function with parameters $a$ and $s$, i.e.,
(4.1) $\widetilde{h}_{1}(t)=\widetilde{h}_{1}(t;a,s)=\frac{1}{\sqrt{a}}h_{1}\left(\frac{t}{a}\right)e^{\pi ist^{2}},\quad a\in{\mathbb{R}}_{+},\ s\in{\mathbb{R}},$
the Zak transform $\mathcal{Z}\widetilde{h}_{1}(z)$ has exactly 3 zeros in $[0,1)^{2}$ and these are determined by (2.32).
As $h_{1}(t)=\widetilde{h}_{1}(t;1,0)$, the crucial point is that no new non-trivial zeros occur in the Zak transform when varying the parameters. This would be comparable to the Gaussian case. Note that Theorem 3.1, Proposition 3.2, and Corollary 3.3 hold for generalized Hermite functions. Thus, Theorem 3.4 can be extended to periodic configurations $\Gamma(\Lambda,\{z_{m}\}_{m=1}^{M})$, $\Lambda\subset{\mathbb{R}}^{2}$ a lattice with $D(\Lambda)=1$, by using the results presented in [7].
For a dilated second Hermite function, we now prove the existence of a Gabor system with an oversampling rate of 5, which is not a frame. As shown in [3] or [21, Lem. 3], for $\omega=1/2$, $\mathcal{Z}(\mathcal{D}^{-1}_{a}h_{2})(x,1/2)$ possesses three zeros for $a\in\{\sqrt{2},\sqrt{3}\}$. We pick $a=\sqrt{2}$ and prove the existence of two more zeros: $z_{0}=(x_{0},0)$ and $\widetilde{z}_{0}=(1-x_{0},0)$ with $x_{0}\in(0,1/2)$.
The second Hermite function is explicitly given by
(4.2) $h_{2}(t)=2^{-1/4}(-1+4\pi t^{2})e^{-\pi t^{2}}.$
The Zak transform $\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(x,0)$ with $\omega=0$ is a real-valued, continuous, even, periodic function on ${\mathbb{R}}$, depicted in Figure 6. By using (2.28) and $\mathcal{F}\mathcal{D}^{-1}_{a}h_{2}=-\mathcal{D}_{a}h_{2}$, we obtain
(4.3) $\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(x,0)$ $=-\mathcal{Z}\left(\mathcal{D}_{\sqrt{2}}h_{2}\right)(0,-x)$
(4.4) $=2^{-1/2}\sum_{k\in{\mathbb{Z}}}(1-2\pi k^{2})e^{-\frac{\pi}{2}k^{2}}e^{-2\pi ikx}$
(4.5) $=2^{-1/2}\sum_{k\in{\mathbb{Z}}}(1-2\pi k^{2})e^{-\frac{\pi}{2}k^{2}}\cos(2\pi kx).$
Figure 6. The Zak transform $\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(x,0)$ on $[0,1]$.
We multiply by $2^{1/2}$ and evaluate at $x=0$ and $x=1/2$. Note that $e^{-\frac{\pi}{2}}>1/5$.
(4.6) $(x=0):\quad 2^{1/2}\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(0,0)=1+2\sum_{n\geq 1}\underbrace{(1-2\pi n^{2})}_{<0}e^{-\frac{\pi}{2}n^{2}}<1+2(1-2\pi)e^{-\frac{\pi}{2}}<0.$
(4.7) $(x=1/2):\quad 2^{1/2}\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(1/2,0)$ $=\sum_{k\in{\mathbb{Z}}}(-1)^{k}(1-2\pi k^{2})e^{-\frac{\pi}{2}k^{2}}$
(4.8) $>1+2(-1+2\pi)e^{-\frac{\pi}{2}}-2\sum_{n\geq 2}(1+2\pi n^{2})e^{-\frac{\pi}{2}n^{2}}$
(4.9) $>1+2(-1+2\pi)e^{-\frac{\pi}{2}}-2\sum_{n\geq 2}(1+2\pi n^{2})e^{-\pi n}.$
We derive closed-form expressions for the series in the Appendix. The analytic expressions can be evaluated to arbitrary precision by hand. However, we used Mathematica [28], which can actually evaluate the series symbolically, to obtain the numerical bound
(4.10) $32\sum_{n\geq 2}(1+2\pi n^{2})e^{-\pi n}<0.11\quad\text{ while }\quad 1+2(-1+2\pi)e^{-\frac{\pi}{2}}>$ 3 .
Thus, it follows that $\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(1/2,0)>0$, which proves the existence of a zero $(x_{0},0)$ of $\mathcal{Z}\left(\mathcal{D}^{-1}_{\sqrt{2}}h_{2}\right)(x,0)$ with $x_{0}\in(0,1/2)$. By symmetry, another zero $(1-x_{0},0)$ exists. Thus, combined with the results from [3] or [21], we know that $\mathcal{D}_{\sqrt{2}}^{-1}h_{2}$ has at least five distinct zeros in $[0,1)^{2}$. This guarantees the existence of a Gabor system with $\mathcal{D}_{\sqrt{2}}^{-1}h_{2}$ and a periodic structure of density 5, which is not a frame. We leave the exact location of the new zeros open. Numerically, they are approximately at $x_{0}=0.171$ and $\widetilde{x}_{0}=0.829$.
Lastly, we note that $\mathcal{Z}_{\sqrt{3}}h_{3}(z)$ (Figure 4) possesses 2 additional zeros for $\omega=\frac{1}{2\sqrt{3}}$ at approximately $x_{0}=0.256\cdot\sqrt{3}$ and $\widetilde{x}_{0}=1-x_{0}=0.744\cdot\sqrt{3}$, as indicated in Figure 7.
Figure 7. The dilated Zak transform $\mathcal{Z}_{\sqrt{3}}h_{3}\left(x,\tfrac{1}{2\sqrt{3}}\right)$ on $[0,\sqrt{3}]$. For $\ell\in{\mathbb{Z}}$, $\mathcal{Z}_{\sqrt{3}}h_{3}\left(x+\ell,\tfrac{1}{2\sqrt{3}}\right)=(-1)^{\ell}\mathcal{Z}_{\sqrt{3}}h_{3}\left(x+\ell,\tfrac{1}{2\sqrt{3}}\right)$ by the quasi-periodicity.
We estimated the location using Mathematica [28]. The proof of their existence is similar to that just above. Thus, there exists a periodic structure $\Gamma=\bigcup_{k=1}^{7}\left(D_{\sqrt{3}}{\mathbb{Z}}^{2}+z_{k}\right)$ which has density 7 such that the Gabor system $\mathcal{G}(h_{3},\Gamma)$ is not a frame.
Appendix
We use the following identities. For $|q|<1$, we have
(4.11) $f_{N}(q)=\sum_{n\geq N}q^{n}$ $=\frac{q^{N}}{1-q},$
(4.12) $qf_{N}^{\prime}(q)=\sum_{n\geq N}nq^{n}$ $=\frac{Nq^{N}}{1-q}+\frac{q^{N+1}}{(1-q)^{2}},$
(4.13) $q^{2}f_{N}^{\prime\prime}(q)=\sum_{n\geq N}n^{2}q^{n}$ $=\frac{N^{2}q^{N}}{1-q}+\frac{Nq^{N+1}}{(1-q)^{2}}+\frac{(N+1)q^{N+1}}{(1-q)^{2}}+2\frac{q^{N+2}}{(1-q)^{3}}$
(4.14) $=q^{N}\frac{N^{2}(1-q)^{2}+(2N+1)q(1-q)+2q^{2}}{(1-q)^{3}}$
(4.15) $=q^{N}\frac{\left(N+(1-N)q\right)^{2}+q}{(1-q)^{3}}.$
Picking specifically $N=2$ and $q=e^{-\pi}$, we obtain
(4.16) $\sum_{n\geq 2}e^{-\pi n}$ $=\frac{e^{-2\pi}}{1-e^{-\pi}}\approx 0.00195179,$
(4.17) $\sum_{n\geq 2}n^{2}e^{-\pi n}$ $=\frac{e^{-2\pi}\left(\left(2-e^{-\pi}\right)^{2}+e^{-\pi}\right)}{\left(1-e^{-\pi}\right)^{3}}\approx 0.0082588.$
Together, this leads to the bound $0.112\sum_{n\geq 2}(1+2\pi n^{2})e^{-\pi n}<$ 0.11.